# Investigating a Model-Agnostic and Imputation-Free Approach for Irregularly-Sampled Multivariate Time-Series Modeling

### v2

@@ Abstract
Modeling Irregularly-sampled and Multivariate Time Series (IMTS) is crucial across a variety of applications where different sets of variates may be missing at different time-steps due to sensor malfunctions or high data acquisition costs. Existing approaches for IMTS either consider a two-stage impute-then-model framework or involve specialized architectures specific to a particular model and task. We perform a series of experiments to derive novel insights about the performance of IMTS methods on a variety of semi-synthetic and real-world datasets for both classification and forecasting. We also introduce Missing Feature-aware Time Series Modeling (MissTSM) or MissTSM, a novel model-agnostic and imputation-free approach for IMTS modeling. We show that MissTSM shows competitive performance compared to other IMTS approaches, especially when the amount of missing values is large and the data lacks simplistic periodic structures—conditions common to real-world IMTS applications.
@@ 1 Introduction
Deep learning for modeling Multivariate Time-Series (MTS) is a rapidly growing field in Artificial Intelligence (AI) especially with the use of transformer architectures for capturing temporal dependencies via attention mechanisms (Dong et al. 2024; Nie et al. 2022a; Liu et al. 2023). A common feature of these architectures is to treat a fixed set of variates at every time-step or a fixed set of time-steps for every variate as a token represented using vector embeddings. Research in this field has been especially fueled by the availability of benchmark MTS datasets spanning diverse applications such as electric load forecasting and health monitoring containing fixed sets of variates regularly sampled over time.
However, real-world MTS applications are plagued by missing values occurring over arbitrary sets of variates at every time-step (e.g., due to sensor malfunctions), resulting in Irregularly-sampled MTS (IMTS) datasets. IMTS modeling is particularly challenging because the misalignment of variates across time impairs transformer models that assume a fixed set of variates to be observed at every time-step, unless specific adaptations are made (Shukla and Marlin 2020).
A common approach for IMTS modeling is to use a two-step framework where we first use imputation methods (Ahn, Sun, and Kim 2022; Batista, Monard et al. 2002) to fill in missing values based on observed data, followed by feeding the imputed time-series to an MTS model (see Figure 1). Note that the choice of imputation method is agnostic to the MTS model, making it “model-agnostic.” However, the effectiveness of this framework relies on the quality of performed imputation, which can degrade if the time-series lacks periodic structure or if the imputation method is overly simplistic. Imputation can also introduce artificial patterns or artifacts into the data, which MTS models may interpret as genuine trends or observations.
@@
Imputation-free approaches for IMTS have also been developed in recent literature, including GRU-D (Che et al. 2018) and Latent ODE (Rubanova, Chen, and Duvenaud 2019), that involve specialized architectures to handle missing values in time-series for specific downstream tasks such as classification (see Figure 1). However, these approaches struggle with capturing long-term temporal dependencies that are central to the problem of forecasting, are difficult to parallelize, and incur high computational costs. Furthermore, their architectural innovations for handling missing values are designed for specific tasks and cannot be transferred across MTS models.
Given the complementary strengths and weaknesses of existing approaches in IMTS modeling, we investigate the landscape of research in this field through the lens of the following questions: (1) can we develop a model-agnostic and imputation-free approach for IMTS modeling that can be used in a variety of downstream tasks (e.g., forecasting and classification)? (2) how sensitive are imputation-based frameworks on the choice of imputation technique and the nature of missing values in the data? and (3) how do IMTS approaches fare as we vary the fraction of missing values in the data?
To answer these questions, we perform a series of experiments to derive novel insights about the performance of IMTS methods on a variety of datasets for both classification and forecasting tasks with synthetic masking techniques as well as real-world occurrence of missing values. We also introduce MissTSM, a novel model-agnostic and imputation-free approach for IMTS modeling, designed as a “plug-and-play” layer that can be integrated into any backbone MTS model to enable handling of missing values without requiring explicit imputation. We show that MissTSM shows competitive performance compared to other IMTS approaches, especially when the amount of missing values is large and the data lacks simplistic periodic structures—conditions common to real-world IMTS applications.
@@ 2 Related Works
Time-series Forecasting. With the introduction of attention mechanisms via transformer models (Vaswani et al. 2017), a number of transformer-based time-series models have been developed in the last few years (Wu et al. 2021), (Nie et al. 2022b), (Dong et al. 2024), (Liu et al. 2023). While Transformer-based models have shown great promise, recently there has been a strong interest in exploring the use of simple linear models for time-series forecasting as well (Zeng et al. 2023; Ekambaram et al. 2023). In addition, with the rise of self-supervised learning-based models such as masked auto-encoders (MAEs) (He et al. 2022), a new category of MAE-style time-series models have emerged (Dong et al. 2024) that have received a lot of recent interest owing to their ability of learning both low-level and high-level representations for varied downstream tasks such as forecasting and classification. However, while these methods can deal with missing values in the temporal domain, they are unable to handle missing values across both variates and time.
@@
Imputation Methods. Traditionally, most imputation techniques for handling missing values in time-series have been based on statistical approaches (Fung 2006; Batista, Monard et al. 2002; Dempster, Laird, and Rubin 1977; Mnih and Salakhutdinov 2007). In recent years, there is a growing trend to use deep learning methods for time-series imputation, such as SAITS (Du, Côté, and Liu 2023), CSDI (Tashiro et al. 2021a), GAIN (Yoon, Jordon, and Schaar 2018), and BRITS (Cao et al. 2018). Imputation techniques can be broadly classified into two classes: those that leverage cross-channel correlations (Batista, Monard et al. 2002; Acuna and Rodriguez 2004) and those that exploit temporal dynamics (Box et al. 2015). Recently, deep learning-based approaches for imputation have been developed (Tashiro et al. 2021b; Cini, Marisca, and Alippi 2021; Liu et al. 2019; Cao et al. 2018; Du, Côté, and Liu 2023), which can jointly learn the temporal dynamics with cross-channel correlations. These methods, however, rely on a single entangled representation (or hidden state) to model nonlinear dynamics (Woo et al. 2022) which can be a limitation in capturing the multifaceted nature of time-series. Matrix factorization based techniques (Liu et al. 2022) have also been proposed that offer disentangled temporal representations, enhancing the ability to differentiate and model distinct temporal features. While these deep learning-based models are highly efficient during inference, they require additional training time, which add to the already large time complexity of MTS models.
@@
Imputation-free IMTS Models. In the last decade, there has been a significant growth of models and architectures for learning from IMTS data. Some of the simpler approaches to deal with IMTS data involve working with fixed temporal discretization (Marlin et al. 2012; Lipton, Kale, and Wetzel 2016). The primary drawback with these approaches is that they make ad-hoc choices in terms of discretization window width and aggregation functions within the windows (Shukla and Marlin 2020). A popular set of approaches for handling IMTS data are recurrence-based approaches, which includes RNN-based methods such as GRU-D (Che et al. 2018). However, GRU-D has limited scalability to long sequences. Other recurrence-based approaches based on Ordinary Differential Equations (ODE) (Chen et al. 2018; Rubanova, Chen, and Duvenaud 2019) provide an effective solution in modeling the continuous time semantics. These methods are however significantly slow and memory-intensive, as they constantly need to apply the ODE solver and solving ODEs require numerical integration, thus making it impractical for long-term forecasting and large datasets.
@@
Transformer-based methods such as ContiFormer (Chen et al. 2023) and mTAN (Shukla and Marlin 2021) addresses these limitations by explicitly integrating the modeling abilities of Neural ODEs into the attention mechanism and introducing continuous time attention mechanism that learns time embeddings dynamically, respectively. However, despite ContiFormer being a principled and effective approach, solving an ODE for each key and value incurs a high computational cost. Also, while the runtime speed of mTAN is relatively faster, it is however, inherently optimized toward interpolating missing values by learning representations at fixed set of reference points, thus limiting it’s extrapolation or forecasting ability.
@@
This is another limitation of IMTS approaches—their evaluation is mostly limited to a single task, most often to time series classification, thus limiting their applicability. ContiFormer performs evaluation on forecasting tasks, however, they consider regular and clean benchmark time-series datasets in their evaluation. Another limitation in terms of evaluation is that the prior works primarily focus on other IMTS models for comparison, completely ignoring the two-stage imputation approach, which is a more common and practical way of dealing with missing-value data. Our work aims to solve these issues by providing a comprehensive comparison against both existing imputation-free and two-stage imputation-based approaches, and proposing a model agnostic transformation-allowing any task-specific SOTA model to be applied on any irregularly sampled time-series data with minimal data transformations.
@@ 3.2 Learning Embeddings for Time-Series with Missing Features
Limitations of Existing Transformer Methods: The first step in time-series modeling using transformer-based architectures is to learn an embedding of the time-series $\mathbf{X}$ that can be sent to the transformer encoder.Traditionally, this is done using an Embedding-layer (typically implemented using a multi-layered perceptron) as $\texttt{Embedding}:\mathbb{R}^{N}\mapsto\mathbb{R}^{D}$ that maps $\mathbf{X}\in\mathbb{R}^{T\times N}$ to the embedding $\mathbf{H}\in\mathbb{R}^{T\times D}$, where $D$ is the embedding dimension. The Embedding layer operates on every time-step independently such that the set of variates observed at time-step $t$, $\mathbf{X}_{(t,:)}$, is considered as a single token and mapped to the embedding vector $\mathbf{h}_{t}\in\mathbb{R}^{D}$ as $\mathbf{h}_{t}=\texttt{Embedding}(\mathbf{X}_{(t,:)})$ (see Figure 2(a)). An alternate embedding scheme was recently introduced in the framework of inverted Transformer (iTransformer) (Liu et al. 2023), where the uni-variate time-series for the $d$-th variate, $\mathbf{X}_{(:,d)}$, is considered as a single token and mapped to the embedding vector: $\mathbf{h}_{d}=\texttt{Embedding}(\mathbf{X}_{(:,d)})$ (see Figure 2(b)).
Figure 2: Schematic of the Time-Feature Independent (TFI) Embedding of MissTSM that learns a different embedding for every combination of time-step and variate, in contrast to the time-only embeddings of Transformer (Vaswani et al. 2017) and the variate-only embeddings of iTransformers (Liu et al. 2023).
@@
Time-Feature Independent (TFI) Embedding: To address this challenge as well as to utilize inter-variate interactions similar to (Wei et al. 2023), we consider a Time-Feature Independent (TFI) Embedding scheme for time-series with missing features, where the value at each combination of time-step $t$ and variate $d$ is considered as a single token $\mathbf{X}_{(t,d)}$, and is independently mapped to an embedding using $\texttt{TFIEmbedding}:\mathbb{R}\mapsto\mathbb{R}^{D}$ as follows:
@@
In other words, the TFIEmbedding Layer maps $\mathbf{X}\in\mathbb{R}^{T\times N}$ into the TFI embedding $\mathbf{H}^{\mathrm{TFI}}\in\mathbb{R}^{T\times N\times D}$ (see Figure 2(c)). The TFIEmbedding is applied only on tokens $\mathbf{X}_{(t,d)}$ that are observed (for missing tokens, i.e., $\mathcal{M}_{(t,d)}=1$, we generate a dummy embedding that gets masked out in the MFAA layer). The advantage of such an approach is that even if a particular value in the time-series is missing, other observed values in the time-series can be embedded “independently” without being affected by the missing values. Moreover, it allows the MFAA layer to leverage the high-dimensional embeddings to store richer representations bringing in the context of time and variate by computing masked cross-attention among the observed features at a time-step to account for the missing features.
@@
Figure 3: Overview of the MissTSM layer integrated within the Masked Auto-Encoder framework (Li et al. 2023). A zoomed-in view of the MFAA is shown on the left.

### v3

@@ Abstract
Modeling Irregularly-sampled and Multivariate Time Series (IMTS) is crucial across a variety of applications where different sets of variates may be missing at different time-steps due to sensor malfunctions or high data acquisition costs. Existing approaches for IMTS either consider a two-stage impute-then-model framework or involve specialized architectures specific to a particular model and task. We perform a series of experiments to derive insights about the performance of IMTS methods on a variety of semi-synthetic and real-world datasets for both classification and forecasting. We also introduce Missing Feature-aware Time Series Modeling (MissTSM) or MissTSM, a simple model-agnostic and imputation-free approach for IMTS modeling. We show that MissTSM shows competitive performance compared to other IMTS approaches, especially when the amount of missing values is large and the data lacks simplistic periodic structures–conditions common to real-world IMTS applications.
footnotetext: *Corresponding author: abhilash22@vt.edu, karpatne@vt.edufootnotetext: Code available at: https://github.com/abhilash-neog/SparseTimeSeriesModeling
@@ 1 Introduction
Deep Learning for modeling multivariate Time-Series (MTS) is a rapidly growing field, with two major downstream tasks: forecasting and classification. Research (Dong et al., 2024; Nie et al., 2022a; Liu et al., 2023) in this field has been fueled by the availability of benchmark MTS datasets spanning diverse applications such as electric load forecasting and health monitoring containing fixed sets of variates regularly sampled over time. However, real-world MTS applications are plagued by missing values occurring over arbitrary sets of variates at every time-step (e.g. due to sensor malfunctions), resulting in Irregularly-sampled MTS (IMTS) datasets. IMTS modeling is particularly challenging because the misalignment of variates across time impairs transformer models that assume a fixed set of variates to be observed at every time-step
A common approach for IMTS modeling is to use a two-step framework where we first use imputation methods (Ahn et al., 2022; Batista et al., 2002) to fill in missing values based on observed data, followed by feeding the imputed time-series to an MTS model (see Figure 1). Note that the choice of imputation method is agnostic to the MTS model, making it “model-agnostic.” However, the effectiveness of this framework relies on the quality of performed imputation, which can degrade if the time-series lacks periodic structure or if the imputation method is overly simplistic. Imputation can also introduce artificial patterns or artifacts into the data, which MTS models may interpret as genuine trends or observations. Moreover, deep learning-based imputation methods require training, which adds to the overall computational cost of IMTS modeling.
Imputation-free approaches for IMTS have also been developed in recent literature (Che et al., 2018; Rubanova et al., 2019), that involve specialized architectures to handle missing values in time-series for specific downstream tasks such as classification (see Figure 1) However, these approaches have been empirically shown to struggle with capturing long-term temporal dependencies that are central to the problem of forecasting, are difficult to parallelize, and incur high computational costs. Furthermore, existing imputation-free IMTS approaches are not model-agnostic, i.e., they have been developed as specialized model architectures that cannot be used as a generic wrapper with latest advances in MTS models, restricting their adaptability and performance.
@@
Given the complementary strengths and weaknesses of existing approaches in IMTS modeling, we ask the following question - can we develop a model-agnostic and imputation-free approach for IMTS modeling that can be used in a variety of downstream tasks (e.g., forecasting and classification)? To address this question, we analyze the prevailing strategies for converting time-series into tokens in existing transformer-based MTS models. There are two primary ways for converting time series into tokens: (a) treating each time step (all variates) as a token, or (b) considering each variate (all time steps) as a token (Liu et al., 2023). While these strategies work well for regular MTS data, they are not suited to handle IMTS data because we may be missing some variates at a time step or some time-steps for a variate, making the computation of tokens infeasible. In contrast, we explore a different perspective for creating time-series tokens: independently embedding each combination of time-step and variate as a token. Time-variate combinations with missing values can then be handled using masked cross-attention without performing any explicit imputation. Building on this intuition, we introduce MissTSM, a simple model-agnostic and imputation-free approach for IMTS modeling, designed as a “plug-and-play” layer that can be integrated into any backbone MTS model to handle IMTS data. The advantage of such an approach is that it (a) does not introduce any imputation artifacts, and (b) can act as a wrapper around any MTS model.
In this work, we make the following contributions: (1) We introduce MissTSM, a model-agnostic and imputation-free approach; (2) We conduct a comprehensive experimental study on a variety of datasets for both classification and forecasting tasks, using synthetic masking techniques as well as real-world occurrence of missing values. This study investigates (a) sensitivity of imputation-based frameworks on the choice of imputation technique and the nature of missing values, and (b) the performance of IMTS approaches as the fraction of missing values varies; (3) We demonstrate that MissTSM achieves competitive performance compared to other IMTS approaches, especially when the amount of missing values is large and the data lacks simplistic periodic structures - conditions common to real-world IMTS applications.
@@ 2 Related Works
Time-series Forecasting. With the introduction of attention mechanisms via transformer models (Vaswani et al., 2017), a number of transformer-based time-series models have been developed in the last few years (Wu et al., 2021; Nie et al., 2022b; Dong et al., 2024; Liu et al., 2023). While Transformer-based models have shown great promise, recently there has been a strong interest in exploring the use of simple linear models for time-series forecasting as well (Zeng et al., 2023; Ekambaram et al., 2023). In addition, with the rise of self-supervised learning-based models such as masked auto-encoders (MAEs) (He et al., 2022), a new category of MAE-style time-series models have emerged (Dong et al., 2024) that have received a lot of recent interest owing to their ability of learning both low-level and high-level representations for varied downstream tasks such as forecasting and classification. However, while these methods can deal with missing values in the temporal domain, they are unable to handle missing values across both variates and time.
@@
Imputation Methods. Traditionally, most imputation techniques for handling missing values in time-series have been based on statistical approaches (Fung, 2006; Batista et al., 2002; Dempster et al., 1977; Mnih and Salakhutdinov, 2007). In recent years, there is a growing trend to use deep learning methods for time-series imputation, such as SAITS (Du et al., 2023), CSDI (Tashiro et al., 2021a), GAIN (Yoon et al., 2018a), and BRITS (Cao et al., 2018). Imputation techniques can be broadly classified into two classes: those that leverage cross-channel correlations (Batista et al., 2002; Acuna and Rodriguez, 2004) and those that exploit temporal dynamics (Box et al., 2015). Recently, deep learning-based approaches for imputation have been developed (Tashiro et al., 2021b; Cini et al., 2021; Liu et al., 2019; Cao et al., 2018; Du et al., 2023), which can jointly learn the temporal dynamics with cross-channel correlations. These methods, however, rely on a single entangled representation (or hidden state) to model nonlinear dynamics (Woo et al., 2022) which can be a limitation in capturing the multifaceted nature of time-series. Matrix factorization based techniques (Liu et al., 2022) have also been proposed that offer disentangled temporal representations, enhancing the ability to differentiate and model distinct temporal features. While these deep learning-based models are highly efficient during inference, they require additional training time, which add to the already large time complexity of MTS models.
@@
Imputation-free IMTS Models. In the last decade, there has been a significant growth of models and architectures for learning from IMTS data. Some of the simpler approaches to deal with IMTS data involve working with fixed temporal discretization (Marlin et al., 2012; Lipton et al., 2016). The primary drawback with these approaches is that they make ad-hoc choices in terms of discretization window width and aggregation functions within the windows (Shukla and Marlin, 2020). A popular set of approaches for handling IMTS data are recurrence-based approaches, which includes RNN-based methods such as GRU-D (Che et al., 2018). However, GRU-D has limited scalability to long sequences. Other recurrence-based approaches based on Ordinary Differential Equations (ODE) (Chen et al., 2018; Rubanova et al., 2019) provide an effective solution in modeling the continuous time semantics. These methods are however significantly slow and memory-intensive, as they constantly need to apply the ODE solver and solving ODEs require numerical integration, thus making it impractical for long-term forecasting and large datasets.
@@
Transformer-based methods such as ContiFormer (Chen et al., 2023) and mTAN (Shukla and Marlin, 2021) addresses these limitations by explicitly integrating the modeling abilities of Neural ODEs into the attention mechanism and introducing continuous time attention mechanism that learns time embeddings dynamically, respectively. However, despite ContiFormer being a principled and effective approach, solving an ODE for each key and value incurs a high computational cost. Also, while the runtime speed of mTAN is relatively faster, it is however, inherently optimized toward interpolating missing values by learning representations at fixed set of reference points, thus limiting it’s extrapolation or forecasting ability.
@@
This is another limitation of IMTS approaches—their evaluation is mostly limited to a single task, most often to time series classification, thus limiting their applicability. ContiFormer performs evaluation on forecasting tasks, however, they consider regular and clean benchmark time-series datasets in their evaluation. Another limitation in terms of evaluation is that the prior works primarily focus on other IMTS models for comparison, completely ignoring the two-stage imputation approach, which is a more common and practical way of dealing with missing-value data. Our work aims to solve these issues by providing a comprehensive comparison against both existing imputation-free and two-stage imputation-based approaches, and proposing a model-agnostic transformation-allowing any task-specific SOTA model to be applied on any irregularly sampled time-series data with minimal data transformations.
@@ 3.2 Learning Embeddings for Time-Series with Missing Features
Limitations of Existing Transformer Methods: The first step in time-series modeling using transformer-based architectures is to learn an embedding of the time-series $\mathbf{X}$ that can be sent to the transformer encoder. Traditionally, this is done using an Embedding layer (typically implemented using a multi-layered perceptron) as $\texttt{Embedding}:\mathbb{R}^{N}\mapsto\mathbb{R}^{D}$ that maps $\mathbf{X}\in\mathbb{R}^{T\times N}$ to the embedding $\mathbf{H}\in\mathbb{R}^{T\times D}$, where $D$ is the embedding dimension. The Embedding layer operates on every time-step independently such that the set of variates observed at time-step $t$, $\mathbf{X}_{(t,:)}$, is considered as a single token and mapped to the embedding vector $\mathbf{h}_{t}\in\mathbb{R}^{D}$ as $\mathbf{h}_{t}=\texttt{Embedding}(\mathbf{X}_{(t,:)})$ (see Figure 2(a)).
Figure 2: Schematic of the Time-Feature Independent (TFI) Embedding of MissTSM that learns a different embedding for every combination of time-step and variate, in contrast to the time-only embeddings of Transformer (Vaswani et al., 2017) and the variate-only embeddings of iTransformers (Liu et al., 2023).
An alternate embedding scheme was recently introduced in the framework of inverted Transformer (iTransformer) (Liu et al., 2023), where the uni-variate time-series for the $d$-th variate, $\mathbf{X}_{(:,d)}$, is considered as a single token and mapped to the embedding vector: $\mathbf{h}_{d}=\texttt{Embedding}(\mathbf{X}_{(:,d)})$ (see Figure 2(b)).
@@
Time-Feature Independent (TFI) Embedding: To address this challenge as well as to utilize inter-variate interactions similar to Wei et al. (2023), we consider a Time-Feature Independent (TFI) Embedding scheme for time-series with missing features, where the value at each combination of time-step $t$ and variate $d$ is considered as a single token $\mathbf{X}_{(t,d)}$, and is independently mapped to an embedding using $\texttt{TFIEmbedding}:\mathbb{R}\mapsto\mathbb{R}^{D}$ as follows:
@@
In other words, the TFIEmbedding Layer (which is a simple MLP layer) maps $\mathbf{X}\in\mathbb{R}^{T\times N}$ into the TFI embedding $\mathbf{H}^{\mathrm{TFI}}\in\mathbb{R}^{T\times N\times D}$ (see Figure 2(c)). The TFIEmbedding is applied only on tokens $\mathbf{X}_{(t,d)}$ that are observed (for missing tokens, i.e., $\mathcal{M}_{(t,d)}=1$, we generate a dummy embedding that gets masked out in the MFAA layer). The advantage of such an approach is that even if a particular value in the time-series is missing, other observed values in the time-series can be embedded “independently” without being affected by the missing values. Moreover, it allows the MFAA layer to leverage the high-dimensional embeddings to store richer representations bringing in the context of time and variate by computing masked cross-attention among the observed features at a time-step to account for the missing features.
@@
Figure 3: Overview of the MissTSM layer integrated within the Masked Auto-Encoder framework (Li et al., 2023). A zoomed-in view of the MFAA is shown on the left.
3.3 Missing Feature-Aware Attention (MFAA)
The MFAA Layer illustrated in Figure 3 leverage the power of “masked-attention” for learning latent representations at every time-step using partially observed features. MFAA works by computing attention scores based on the partially observed features at a time-step $t$, which are then used to perform a weighted sum of observed features to obtain the latent representation $\mathbf{L}_{t}$. As shown in Figure 3, these latent representations are projected back using a linear layer, to the original input shape before being fed into the downstream model (here, the encoder-decoder based self-supervised learning framework). Note that MFAA is not designed for long-term temporal modeling. Its primary role is to handle missing values at a time-step by using the observed variates at the same time-step as context. It relies on the subsequent backbone to model long-range temporal dynamics MFAA performs a masked cross-attention using a learnable query vector and observed data as keys and values. This separation of roles is inspired by similar architectures in multi-modal grounding, for example, in Carion et al. (2020), where learnable object queries serve as abstract object representations to focus on distinct objects in an image without requiring predefined region proposals, enabling set-based prediction. Similarly, in our setting, the learnable queries capture the interactions among variates independent of time, enabling the model to attend to the most informative aspects of observed variates at any time-step fed through keys and values. This intuition aligns with the query-based mechanism in mTAN (Shukla and Marlin, 2021), which introduces a structured way to aggregate information over observed time-series data. However, there’s is a key difference in the nature of the query - while mTAN uses discrete reference points on a fixed temporal grid to achieve this, our single learnable query generalizes across variates at every time step, allowing for a more flexible representation of feature interactions. Another key difference is in the architecture flow - mTAN first performs a "temporal-first, variate-independent interpolation" to perform temporal attention for each feature to create a regular sequence, then performs a linear combination to generate a single per timestep representation. In contrast, MissTSM is "variate-first", with the MFAA layer performing cross-variate attention to create a representation $L_{t}$ at each timestep. Then, the backbone model learns the temporal relationships from this sequence of summaries. Also note that our goal is not to propose a new attention mechanism, rather, our novelty lies in the novel adaptation of masked cross-attention for effectively modeling IMTS data. To the best of our knowledge, this novel adaptation of masked cross-attention for IMTS modeling has not been tried before in any previous works.
Mathematical Formulation: To obtain attention scores from partially observed features at a time-step, we apply a masked scaled-dot product operation followed by a softmax operation described as follows. We first define a learnable query vector $\mathbf{Q}\in\mathbb{R}^{1\times D}$ which is independent of the variates and time-steps. The positionally-encoded embeddings at time-step $t$, $\mathbf{Z}_{(t,:)}$, are used as key and value inputs in the MFAA Layer. Specifically, The query, key, and value vectors are defined using linear projections as follows: $\hat{\mathbf{Q}}=\mathbf{Q}\mathbf{W^{Q}},\quad\hat{\mathbf{K}}_{t}=\mathbf{Z}_{(t,:)}\mathbf{W^{K}},\quad\hat{\mathbf{V}}_{t}=\mathbf{Z}_{(t,:)}\mathbf{W^{V}}$. Here, $\hat{\mathbf{Q}}\in\mathbb{R}^{1\times d_{k}}$ and $\hat{\mathbf{K}}_{t},\hat{\mathbf{V}}_{t}\in\mathbb{R}^{N\times d_{k}}$, where $d_{k}$ is the dimension of the vectors after linear projection. The linear projection matrices for the query, key, and values are defined as: $\mathbf{W^{Q}},\mathbf{W^{K}},\mathbf{W^{V}}\in\mathbb{R}^{D\times d_{k}}$ respectively. Note that the key $\hat{\mathbf{K}}_{t}$ and value $\hat{\mathbf{V}}_{t}$ vectors depend on the time-step $t$, while the query vector doesn’t change with time. We then define the Missing Feature-Aware Attention Score at a given time-step $t$ as a masked scalar dot-product of the query and key vector followed by normalization of the scores using a Softmax operation, formally defined as follows:
where $\mathbf{A}_{t}\in\mathbb{R}^{N}$ is the MFAA Score vector of size $N$ corresponding to the $N$ variates, and $\eta\rightarrow$-$\infty$ is a large negative bias. The negative bias term $\eta$ forces the masked-elements that correspond to the missing variates in the time-series to have an attention score of zero. Thus, by definition, the $i$-th element of the MFAA Score $\mathbf{A}_{(t,i)}\neq 0\implies\mathcal{M}_{(t,:)}=0$. We compute the latent representation $\mathbf{L}_{t}$ as a weighted sum of the MFAA score $\mathbf{A}_{t}$ and the Value vector $\hat{\mathbf{V}}_{t}$ as follows:
$=\mathrm{Concat}(\mathbf{L}^{0}_{t},\mathbf{L}^{1}_{t},\dots,\mathbf{L}^{h-1}_{t})\cdot\mathbf{W}^{O}$
3.4 Putting Everything Together: Plugging MissTSM with any MTS Model
Figure 3 shows the overall framework of a Masked Auto-Encoder (MAE) (He et al., 2022) based time-series model integrated with MissTSM. For an input time-series $\mathbf{X}$, we apply the TFI embedding layer followed by the MFAA layer to learn a latent representation for every time-step. The latent representations are then projected back to the original input shape to be fed into the downstream model. In this work, we opted for a MAE-based time-series model as the default downstream or base model, primarily due to its recent success in time-series modeling and its ability to perform both time-series forecasting and classification tasks. Furthermore, out of the several state-of-the-art masked time-series modeling techniques, we intentionally chose the simplest variation of MAE, namely Ti-MAE (Li et al., 2023), to highlight the effectiveness of TFI and MFAA layers in handling missing values.
4 Experimental Setup
Baselines: We benchmark against two categories of models. For MTS, we consider SimMTM (Dong et al., 2024), PatchTST (Nie et al., 2022b), AutoFormer (Wu et al., 2021), DLinear (Zeng et al., 2023), and iTransformer (Liu et al., 2023). Imputation strategies used are, $2^{\text{nd}}$-order spline interpolation (McKinley and Levine, 1998), k-Nearest Neighbor (Tan et al., 2019), and SAITS (Du et al., 2023) and BRITS (Cao et al., 2018). For IMTS, we evaluate GRU-D (Che et al., 2018), Latent ODE (Rubanova et al., 2019), SeFT (Horn et al., 2020), mTAND (Shukla and Marlin, 2021), Raindrop (Zhang et al., 2021), and MTGNN (Wu et al., 2020). Baseline choice is aligned with the task each model was originally designed for.
Datasets: We considered three popular time-series forecasting datasets: ETTh2, ETTm2 (Zhou et al., 2021) and Weather (Weather, 2021). For classification, we considered three real-world datasets, namely, Epilepsy (Andrzejak et al., 2001), EMG (Goldberger et al., 2000a), and Gesture (Liu et al., 2009). We follow the same evaluation setups as proposed in TF-C (Zhang et al., 2022). To simulate varying scenarios of missing values appearing in real-world time-series datasets, we adopt two synthetic masking schemes that we apply on these benchmark datasets, namely missing completely at random (MCAR) masking and periodic masking. Furthermore, we compared our performance on five real-world datasets: PhysioNet-2012 (Silva et al., 2012), P12 (Goldberger et al., 2000b) and P19 (Reyna et al., 2020) for health monitoring; Falling Creek Reservoir (FCR) dataset for modeling lake water quality, and Lake Mendota from the North Temperate Lakes Long-Term Ecological Research program (NTL-LTER; Magnuson et al., 2024) also for modeling lakes. See Appendix for more details.
5 Results and Discussions
5.1 Imputation-based vs. Imputation-free
5.1.1 Impact of Missing Data Fractions.